OpenStudy (anonymous):

prove that sqrt(5) is irrational :)

OpenStudy (anonymous):

it's not a perfect square, so it's irrational

OpenStudy (anonymous):

how to prove this?

OpenStudy (anonymous):

all square roots of numbers other than perfect squares yields an irrational response, because that's the definition of a perfect square

OpenStudy (anonymous):

i know this already. but if you were given this question to work out. how would you do it?

OpenStudy (asnaseer):

One way is to prove by contradiction. i.e. start off by assuming that \(\sqrt{5}\) IS rational and show that this leads to a contradiction - thus proving that it is irrational.

OpenStudy (anonymous):

if i do this assumption. i can say that \[\sqrt{5} = \frac{ a }{ b }\] then \[5 = \frac{ a ^{2} }{ b ^{2} }\] \[5b ^{2} = a ^{2}\] and then i'm stuck.

OpenStudy (asnaseer):

all steps are correct so far :)

OpenStudy (asnaseer):

you also need to state at the start that a and b are co-prime, i.e. they have no common factors so that a/b is a fraction in its simplest form. now think about what the last step means - what does it say about \(a^2\)

OpenStudy (asnaseer):

yo found:\[a^2=5b^2\]therefore \(a^2\) must be a multiple of ???

OpenStudy (anonymous):

multiple of 5

OpenStudy (asnaseer):

correct. now 5 is a prime number, so if \(a^2\) is a multiple of 5, then a must also be a multiple of 5 - agreed?

OpenStudy (anonymous):

yes...

OpenStudy (asnaseer):

therefore we can express a as:\[a=5m\]where m is some other integer. therefore:\[a^2=(5m)^2=25m^2\]now substitute this back into your expression that relates a^2 to b^2 and see what that leads to.

OpenStudy (anonymous):

i get \[25m ^{2} = 5b ^{2}\]

OpenStudy (asnaseer):

yup - now divide both sides by 5 and see what you can say about \(b^2\)...

OpenStudy (anonymous):

\[5m ^{2} = b ^{2}\]

OpenStudy (asnaseer):

... so \(b^2\) must be a multiple of ???

OpenStudy (anonymous):

or 5, just like \[a ^{2}\] was

OpenStudy (anonymous):

of**

OpenStudy (asnaseer):

yes - and, following the same reasoning as we did for \(a^2\), we conclude that b is also a multiple of 5.

OpenStudy (asnaseer):

so we have concluded that both a and b are multiples of 5. which means they are NOT co-prime, i.e. they have a common factor of 5. this contradicts our original assumptions, therefore \(\sqrt{5}\) must be irrational.

OpenStudy (asnaseer):

I hope that makes sense?

OpenStudy (anonymous):

makes sense if we can say that a rational number is one which when expressed as a fraction has no coprime (except 1)

OpenStudy (asnaseer):

yes - that is how the proof is supposed to start off as - remember I stated it above as: "you also need to state at the start that a and b are co-prime, i.e. they have no common factors so that a/b is a fraction in its simplest form."

OpenStudy (anonymous):

makes sense. thanks! goodnight!

OpenStudy (asnaseer):

yw :) and goodnight to you too.